My AD was working fine but no LED, so I realized I had a couple transistors flipped and that fixed it. However, now bypass just terminates the signal while delay on works fine. How can I troubleshoot the bypass on the stomp? I read somewhere that there may have been a bad batch of em.
Check continuity between lugs with your multimeter. With the lugs flat horizontally, the three columns form three switches. In one position, the middles will be connected to the tops, and in the other, the middles will be connected to the bottoms.
However, it sounds like you may have something connected wrong, possibly simply swapped a couple wires. Pics?